OpenStudy (anonymous):

logarithm. solve the equation. log(x)(1/1296)=4 my answer choices are a) 1/6 b) 6 c) 4 d) 1

OpenStudy (anonymous):

is that \[\log_{x}(1/1296) = 4 \]? if so then 1296 = 2^4 * 3 ^4 = 6^4 and therefore x = 6

OpenStudy (anonymous):

Again the rules of logs are that \[\log_x A= n\] gives you \[x^n=A\] In your case you have \(A= 1/1296\) and \(n=4\) giving you \[x^4=\frac{1}{1296}\] Now try your possible choices and see which one when multiplied my itself 4 times gives you 1/1296. (that is answer a)

OpenStudy (anonymous):

Whoops. apologies. 1/1296. HelenMcA is right >.< i gave the answer to 1296.